OpenStudy (anonymous):

Solve for the following equation: -3/(7x)+5/(14x)=1 Solve for the system for x 4x+2y=16 x+y=4 Solve for the system for x 2x+y=24 5x-y=-3 Solve for x 5x= 3-10x Please and Thank you. Also please draw it out for me. I appreciate for time and help.

OpenStudy (whpalmer4):

\[\frac{-3}{7x}+\frac{5}{14x} = 1\]Multiply both sides by 14x to get rid of the fractions. Solve for x. \[4x+2y = 16\]\[x+y=4\]Multiply the second equation by -2, add to the first. Solve for x. \[2x+y=24\]\[5x-y=-3\]Add the two equations. Solve for x. \[5x=3-10x\]Add 10x to both sides. Simplify to solve for x. You need to actually *do* the problems to learn anything...do you have trouble doing them, or do you not understand what to do?

OpenStudy (anonymous):

i understand when it's explained and written out for me. i need to see the steps.

OpenStudy (whpalmer4):

I've described what to do for each of the problems. Try to do what I suggested, and post your work.

OpenStudy (anonymous):

Okay. Thank you.

OpenStudy (anonymous):

I passed my quiz with your help. I appreciate it.

OpenStudy (whpalmer4):

If my help was really doing the job, you wouldn't need my help any longer! :-)

OpenStudy (whpalmer4):

So, what do you get for the first equation after you multiply everything by 14x?

OpenStudy (whpalmer4):

Do you see why I chose 14x, and not say 7x or 9x?

OpenStudy (anonymous):

yes.

OpenStudy (whpalmer4):

Okay, so what do you get for the first equation after you multiply everything by 14x?

OpenStudy (anonymous):

I'm sorry but i had a long day today and need some sleep before I do anymore math.